OpenStudy (itsbribro):

What is the solution to the equation? (x+8)-14=-15 x=

OpenStudy (anonymous):

x+8-14=-15 x-6=-15 x=-9

OpenStudy (x3_drummerchick):

(x +8) - 14 = -15 (x+8) = -1 x= -9

OpenStudy (anonymous):

mathway.com is a good helper with this stuff

OpenStudy (itsbribro):

thank you guys so much! @somethingone @x3_drummerchick

OpenStudy (aq99):

x + 8 - 14= -15 x - 6= -15 x= -9

OpenStudy (aq99):

sorry open study lagged

OpenStudy (itsbribro):

At 1 p.m., Lin records an outside temperature of –14.8°F. At 3 p.m., she records the outside temperature and notices the temperature had increased by 22.6 degrees. What was the temperature that she recorded at 3 p.m.? ______°F

OpenStudy (aq99):

37.4

OpenStudy (itsbribro):

thank you :)

OpenStudy (itsbribro):

wouldnt it be 7.8

OpenStudy (aq99):

no

OpenStudy (itsbribro):

ok

OpenStudy (itsbribro):

i got it wrong it said the answer is 7.8

OpenStudy (aq99):

oops sorry

OpenStudy (itsbribro):

its ok